OpenStudy (anonymous):

Please Help! Medal and Fan!!! Is –10 + (–15) = –15 + (–10) A. commutative property B. associative property C. opposite of a sum property ???

OpenStudy (anonymous):

–10 + (–15) = –15 + (–10)

OpenStudy (anonymous):

a

OpenStudy (anonymous):

the answer would be a

OpenStudy (ilovecake):

a.....

OpenStudy (anonymous):

BOOM! im good!

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

can u give me a medal

OpenStudy (anonymous):

Can you help me with another?

OpenStudy (ilovecake):

no I think the only one who deserves it would be @bradenhart

OpenStudy (anonymous):

sure!!!

OpenStudy (anonymous):

Thanks :) @Ilovecake

OpenStudy (anonymous):

4 + (29 + 16) = 4 + (16 + 29) Is it A. commutative property B. associative property C. opposite of a sum property

OpenStudy (ilovecake):

:) np @bradenhart

OpenStudy (anonymous):

Does anyone know?

OpenStudy (anonymous):

I gave raven a \medal she deserves it:)

OpenStudy (anonymous):

it looks likes c

OpenStudy (ilovecake):

yeah I think it is C too.

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

and thank you @cjackson2002

OpenStudy (anonymous):

1 more?

OpenStudy (ilovecake):

sure

OpenStudy (anonymous):

sure!!!!

OpenStudy (anonymous):

you know I got your back girlie:)

OpenStudy (anonymous):

–12 + (12 + 27) = (–12 + 12) + 27

OpenStudy (anonymous):

i got yours too!!!:)

OpenStudy (anonymous):

–12 + (12 + 27) = (–12 + 12) + 27

OpenStudy (anonymous):

<3

OpenStudy (anonymous):

A. commutative property B. associative property C. opposite of a sum property

OpenStudy (anonymous):

can you give the answer choices

OpenStudy (anonymous):

I think b:) do you do k12 @FDV

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

yea b is the answer

OpenStudy (anonymous):

cool i do k12 tooo

OpenStudy (anonymous):

Thanks! 7th grade

OpenStudy (ilovecake):

same here.

OpenStudy (anonymous):

all 3 of us coolness:)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

i am in 7th grade too!!!:)

OpenStudy (anonymous):

awesome!

OpenStudy (anonymous):

me too:D

OpenStudy (ilovecake):

I am always being left out of everything. :'( I am in k12 and I am in 7th grade. :'(.

OpenStudy (anonymous):

cool @Ilovecake

OpenStudy (ilovecake):

just leave me alone..